OpenStudy (anonymous):

I have a function f(x)=sqrt(2x+1). I need to find point c such that f'(c)=1.....

OpenStudy (anonymous):

So I suppose I have to differentiate (? find the general formula for the derivative?), but I'm not sure how to apply the basic differentiation rules here

OpenStudy (anonymous):

ok so if \[\sqrt{2x+1}=(2x+1)^\frac{ 1 }{ 2 }\] then using differentiation I get \[\frac{ 1 }{ 2 }(2x+1)^\frac{ -1 }{ 2 }\]

OpenStudy (anonymous):

which is \[\frac{ 1 }{ 2 }*\frac{ 1 }{ (2x+1)^\frac{ 1 }{ 2 } }\]

OpenStudy (stamp):

You forgot your chain rule. http://tutorial.math.lamar.edu/Classes/CalcI/ChainRule.aspx Your derivative f'(x) = 1/(2x+1)

OpenStudy (stamp):

You want a point c such that f'(c) = 1 So if 1 = 1/(2c+1) 2c + 1 = 1 2c = 0 c = 0 f'(0) = 1

OpenStudy (stamp):

veification f'(0) = 1 http://www.wolframalpha.com/input/?i=1%2Fsqrt%281%2B2+x%29+%3D+1

OpenStudy (stamp):

\[f(x) = (2x+1)^{1/2}\]\[u=2x+1\]\[du/dx=2\]\[y=u^{1/2}\]\[dy/dx=dy/du*du/dx\]\[y'=\frac{1}{2\sqrt u}2\]\[y'=1/\sqrt{2x+1}\]

OpenStudy (stamp):

The 1/2 and the 2 cancel out

OpenStudy (anonymous):

wait, but how do you get rid of the square root?

OpenStudy (stamp):

I did not rationalize the denominator, the problem was to solve for a point c that yields f'(c) = 1, not to simplify the denominator of my derivative.

OpenStudy (stamp):

If you wanted to:\[f'(x)=\frac{1}{\sqrt{2x+1}}*\frac{\sqrt{2x+1}}{\sqrt{2x+1}}\]\[f'(x)=\frac{\sqrt{2x+1}}{2x+1}\]

OpenStudy (anonymous):

oh so it's just a question of 0 being an obvious solution?

OpenStudy (stamp):

Well we want a point x such that f'(x) = 1\[f'(x)=1/\sqrt{2x+1}\]\[f'(x)=1\]Substitute\[1=1/\sqrt{2x+1}\]Cross multiply\[\sqrt{2x+1}=1\]Square both sides\[2x+1=1^2\]\[2x+1=1\]\[2x=0\]\[x=0/2\]\[x=0\]

OpenStudy (stamp):

So f'(0) = 1

OpenStudy (anonymous):

ok, thanks stamp